RaDyO News and Notes


Hi Tommy and Steve,

The DPOL data from Hawaii is up on my website. The addresses are:

The inwater data is in:
http://www.physics.miami.edu/optics/SkyPolRadiance/Hawaii%20cruise,%20Sep
tember%2009/Intro%20to%20Hawaii%20Data.html

the sky data is in:
http://www.physics.miami.edu/optics/SkyPolRadiance/Hawaii%20cruise,%20Sep
tember%2009/Intro%20to%20Hawaii%20Data.html

We have been checking to make sure the links work, but I am sure we have
missed some, so if you notice something please e-mail me.

I think there are a couple of days left for in-water data...but this is
pretty much it. When this rush passes I am going to go back and get the
same sort of pages up for the Santa Barbara data.

To get to the data files there is a password, which I will give to
whoever asks.....however you can get to the youtube videos without
this....

This website is somewhat rushed, so I will be going back and editing the
english (partly Nepaleze) as time allows.

Ken

--------------------------------------------------------------------------------------02/09/10

FLIP Status Report 160909

Note: Report compiled at 1230 local time.
Position: 21-12.942' N 157-58.276' W
Skies: Clear
Winds: 20 knots from 95 degrees
Seas: 3-4 ft from E direction
Swells: 4-6 ft from NNE direction
Pressure: 30.03 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phones now operational
Comments: FLIP is currently in tow at 6.8 knots just outside of Honolulu and on schedule for its 1300 arrival. Offload of science equipment is scheduled for tomorrow at 0900. A word of praise has to be mentioned for the FLIP crew and MSC who have both worked very hard to support the efforts of the science party over the past several weeks in ensuring a successful overall operation.

Nick

--------------------------------------------------------------------------------------09/16/09

FLIP Status Report 150909b

Note: Updated report compiled at 1030 local time.
Position: 17-54.646' N 159-41.730' W
Skies: Cloudy
Winds: 10 knots from 85 degrees
Seas: 2-3 ft from E direction
Swells: 3-4 ft from NNE direction
Pressure: 29.92 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phones not operational
Comments: FLIP has successfully flipped to a horizontal position and is presently being towed back to Pearl Harbor by Navajo at 8.8 knots. Internet communications have been restored. RTP is still scheduled for Thursday am but the possibility of a return on Wednesday at 1300 has been brought up by the captain of the Navajo who has initiated dialogue with MSC about this option.

Nick

--------------------------------------------------------------------------------------09/15/09

FLIP Status Report 150909a

Note: Report compiled at 0700 local time.
Position: 17-52.549' N 159-42.413' W
Skies: Cloudy
Winds: 10 knots from 85 degrees
Seas: 2-3 ft from E direction
Swells: 3-4 ft from NNE direction
Pressure: 29.92 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 239 nm on an average course of 268 degrees since going vertical. Presently we are preparing to flip horizontal with the USNS Navajo on location as scheduled. The Kilo Moana is standing by until we're secure.

Nick

--------------------------------------------------------------------------------------09/15/09

FLIP Status Report 140909b

Note: Report compiled at 2000 local time.
Position: 17-52.576' N 159-36.303' W
Skies: Overcast
Winds: 12 knots from 80 degrees
Seas: 2-3 ft from E direction
Swells: 3-4 ft from NNE direction
Pressure: 29.91 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 233 nm on an average course of 268 degrees since going vertical. Currently we are on a course of 280 degrees at 0.7 knots. The remaining port and starboard booms were recovered today and the lab space was secured for flipping. USNS Navajo is scheduled to arrive on site tomorrow at 0630 at which time the Kilo Moana will depart. FLIP RTP scheduled for early Thursday morning.

Nick

--------------------------------------------------------------------------------------09/14/09

FLIP Status Report 140909a

Note: Report compiled at 0745 local time.
Position: 17-52.511' N 159-31.446' W
Skies: Clear
Winds: 10 knots from 95 degrees
Seas: 2-3 ft from E direction
Swells: 3-4 ft from NNE direction
Pressure: 29.91 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 228 nm on an average course of 268 degrees since going vertical. Currently we are on a course of 270 degrees at 0.4 knots. All instruments recovered from the booms yesterday as well as the face boom. The other booms will be recovered today and the science lab will be prepared for flipping. Navajo scheduled to arrive on site tomorrow at 0630.

Nick

--------------------------------------------------------------------------------------09/14/09

FLIP Status Report 130909b

Note: Report compiled at 1815 local time. Twice daily reports will be transmitted until tug arrives.
Position: 17-51.898' N 159-25.574' W
Skies: Overcast
Winds: 10 knots from 95 degrees
Seas: 2-3 ft from E direction
Swells: 3-4 ft from NNE direction
Pressure: 29.89 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 223 nm on an average course of 268 degrees since going vertical. Currently we are on a course of 290 degrees at 0.4 knots. Instrument recovery for the booms continued today and all equipment was removed. The face boom was recovered this evening. The remaining booms will be recovered tomorrow and the lab will be secured for flipping.

Nick

--------------------------------------------------------------------------------------09/13/09

FLIP Status Report 130909a

Note: Report compiled at 0830 local time.
Position: 17-50.800' N 159-20.620' W
Skies: Clear
Winds: 10 knots from 90 degrees
Seas: 2-3 ft from E direction
Swells: 3-5 ft from NNE direction
Pressure: 29.94 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 218 nm on an average course of 267 degrees since going vertical. Currently we are on a course of 290 degrees at 0.4 knots. Instrument recovery for the booms continues today. No scheduled deployments of subsurface instruments.Boom recovery begins tomorrow.

Nick

--------------------------------------------------------------------------------------09/13/09

12 September, 2009. KM drifting alongside FLIP, position at 17:00(ZT) was ~ 17 deg 50'N 159 deg 12'W. Winds were ~ 10-15 knots throughout the day, with seas/swells ~ 1-1.5 meters. Skies were mostly clear with occasional cloud.

Deck operations commenced at 0600 with a VMP deployment. An early morning MASCOT drop took place. MASCOT, Acoustic Buoy, RADCAM and Hyperpro were deployed in the morning, a CTD (500 m, with bottles) at noon, and VMP, MASCOT + acoustic resonator, and RADCAM throughout the afternoon. The wave LIDAR operated throughout. A 20 M x 20 M spatial survey is underway throughout the night w/KLUDGE (CTD, Biolume, rad/irrad, AC-9, other stuff) drops to 120 m every 2.5 M.

All science equipment performed nominally, and all planned operations were successfully completed. The plan for Sunday is attached.

NOTE: Science meeting at 18:45.

Regards,

Marlon

Scientific Operations for 13 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/13/09

FLIP Status Report 120909

Note: The internet was not functioning this morning. Report compiled at 1245 local time.
Position: 17-49.226' N 159-11.783' W
Skies: Clear
Winds: 10 knots from 110 degrees
Seas: 2-4 ft from E direction
Swells: 4-5 ft from NNE direction
Pressure: 29.93 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 210 nm on an average course of 267 degrees since going vertical. Currently we are on a course of 290 degrees at 0.5 knots. Instrument recovery for the booms will begin today. Ken Voss' group will be deploying their DPOL and Marlon Lewis' RADCAM from the starboard boom today. Dariusz Stramski's group will be deploying their optical porcupine instrument from the port boom during the day with Ken Melville's group deploying their underwater imaging/turbulence package at night. KM has confirmed their intention of staying on site until the Navajo arrives on location.

Nick

--------------------------------------------------------------------------------------09/12/09

11 September, 2009. KM drifting alongside FLIP, position at 20:00(ZT) was 17 deg 45.2'N 159 deg 01.6'W. Winds were ~ 15-20 knots throughout the day, with seas/swells ~ 1.5-2 meters. Skies were broken cloud throughout.

We completed a 2000 m CTD in the early morning (0100-0300) following completion of an underway spatial survey. Deck operations commenced at 0600 with a VMP deployment. An early morning MASCOT drop took place. MASCOT, Acoustic Buoy, RADCAM and Hyperpro were deployed in the morning, a CTD (500 m, with bottles) at noon, and VMP, MASCOT + acoustic resonator, and RADCAM throughout the afternoon. VMP, MASCOT, acoustic resonator and BIOLUME are deploying this evening leading to an 0100 CTD. The wave LIDAR operated throughout.

All science equipment performed nominally, and all planned operations were successfully completed. The plan for Saturday is attached.

NOTE: Science meeting at 18:45.

Regards,

Marlon

Scientific Operations for 12 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/12/09

FLIP Status Report 110909

Note: Compiled at 0830 local time
Position: 17-43.595' N 158-53.077' W
Skies: Clear
Winds: 20 knots from 110 degrees
Seas: 3-5 ft from E direction
Swells: 5-7 ft from NNE direction
Pressure: 29.91 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 193 nm on an average course of 265 degrees since going vertical. Currently we are on a course of 290 degrees at 1 knot. All instruments for the booms and the crow's nest have been deployed and are running continuously. Ken Voss' group will be deploying their sky camera from the face boom along with their DPOL and Marlon Lewis' RADCAM from the starboard boom today. Nick Hall Patch will be deploying his acoustic resonator off the starboard side as well. Finally, Dariusz Stramski's group will be deploying their underwater imaging/turbulence package today from the port boom during the day with Ken Melville's group deploying their underwater imaging/turbulence package at night. Recovery of instrumentation commences tomorrow.

Nick

--------------------------------------------------------------------------------------09/11/09

10 September, 2009. KM drifting alongside FLIP, position at 17:00(ZT) was 17 deg 39.4'N 158 deg 40.5'W. Winds were ~ 15-22 knots throughout the day, with seas/swells ~ 1.5-2 meters. Skies were mostly overcast with periodic squalls.

Operations commenced at 0600 with a VMP deployment. An early morning MASCOT drop took place. MASCOT,Acoustic Buoy, RADCAM and Hyperpro were deployed in the morning, a CTD (500 m, with bottles) at noon, and VMP, MASCOT + acoustic resonator, and RADCAM throughout the afternoon. We are currently carrying out a spatial survey (surface waters only) over a 20 M box, followed by a 0100 CTD cast to 2000 m (w/o bottles). The wave LIDAR operated throughout.

All science equipment performed nominally, and all planned operations were successfully completed.

NOTE: Science meeting at 18:45.

Regards,

Marlon

Scientific Operations for 11 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/11/09

FLIP Status Report 100909

Note: Compiled at 0930 local time

Position: 17-37.742' N 158-34.070' W

Skies: Overcast

Winds: 18 knots from 65 degrees

Seas: 3-5 ft from E direction

Swells: 5-7 ft from NNE direction

Pressure: 29.87 in and steady

Equipment: All FLIP equipment operational

Fuel consumed last 24 hours: 80 gallons

Communications: Cell phones not operational

Comments: FLIP has now drifted 175 nm on an average course of 263 degrees since going vertical. Yesterday Chris Zappa installed a second sonic anemometer on the starboard boom as a supplemental measurement. Otherwise all instruments for the booms and the crow's nest have now been deployed and are running continuously. Ken Voss' group will be deploying their sky camera from the face boom along with their DPOL and Marlon Lewis' RADCAM from the starboard boom today. Nick Hall Patch will be deploying his acoustic resonator off the starboard side as well. Finally, Ken Melville's group will be deploying their underwater imaging/turbulence package today from the port boom during the day as it is too overcast for the porcuping to operate. Tom Golfinos switched generators yesterday at 1340. Transferred trash with KM small boat yesterday morning as well.

Nick

--------------------------------------------------------------------------------------09/10/09

Marlon

Scientific Operations for 10 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/10/09

FLIP Status Report 090909

Note: Compiled at 0800 local time
Position: 17-36.788' N 158-14.900' W
Skies: Clear
Winds: 15 knots from 95 degrees
Seas: 3-4 ft from E direction
Swells: 4-6 ft from NNE direction
Pressure: 29.86 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 157 nm on an average course of 261 degrees since going vertical. Currently we are on a directly westward heading. All instruments for all booms and the crow's nest have now been deployed and are running continuously. The WAMOS was successfully initiated yesterday evening. Ken Voss' group will be deploying their DPOL and sky camera from the face boom and Marlon Lewis' RADCAM from the starboard boom today. Nick Hall Patch will be deploying his acoustic resonator off the starboard side today. Finally, Dariusz Stramski's group will be deploying their optical porcupine instrument from the port boom during the day with Ken Melville's group deploying their underwater imaging/turbulence package after dark. Small boat operations proceed today from the KM with trash transfer from FLIP.

Nick

--------------------------------------------------------------------------------------09/09/09

8 September, 2009. KM drifting alongside FLIP, position at 20:00(ZT) was 17 deg 33.5'N 158 deg 03.6'W. Winds were ~ 18-25 knots throughout the day, with seas/swells ~ 2-2.5 meters. Skies were variable, but mostly clear.

Operations commenced at 0600 with a VMP deployment. An early morning MASCOT drop took place. Unsettled seas in the morning resulted in cancellation of small boat launch. MASCOT,Acoustic Buoy, RADCAM, and Hyperpro were deployed in the morning, a CTD (500 m, with bottles) at noon, and VMP, MASCOT + acoustic resonator, RADCAM and KLUDGE throughout the afternoon. RADCAM deployments on KM were coordinated with RADCAM and surface wave observations on FLIP in both morning and afternoon. A 0100 CTD> cast (w/o bottles) is planned. The wave LIDAR operated throughout.

All science equipment performed nominally, and all planned operations were successfully completed with the exception of those requiring small boat operations. The plan for tomorrow is attached.

NOTE: Science meeting at 18:45.

Regards,

Marlon

Scientific Operations for 9 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/09/09

FLIP Status Report 080909

Note: Compiled at 0800 local time
Position: 17-33.023' N 157-56.223' W
Skies: Cloudy
Winds: 22 knots from 70 degrees
Seas: 3-5 ft from E direction
Swells: 5-7 ft from NNE direction
Pressure: 29.90 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 140 nm on an average course of 259 degrees since going vertical. Currently we are on more of a direct westerly course. Yesterday Marlon Lewis' albedo sensor was deployed on the port boom. Other than that all instruments for the all booms and the crowe's nest have now been deployed and are running continuously. A failed UPS system prevented deployment of Eric Terrill's WAMOS system yesterday but an effort will be made to deploy the system again today. Ken Voss' group will be deploying their DPOL and sky camera from the face boom and Marlon Lewis' RADCAM from the starboard boom, Nick Hall Patch will be deploying his acoustic resonator from the starboard side, and Ken Melville's group will be deploying their underwater imaging/turbulence package from the port boom. They switched with Dariusz Stramski's group today as the conditions are less favorable for their porcupine instrument.

Nick

--------------------------------------------------------------------------------------09/08/09

7 September, 2009. KM drifting alongside FLIP, position at 20:00(ZT) was 17 deg 33.4'N 157 deg 47.6'W. Winds were ~ 15-20 knots throughout the day, with seas/swells ~ 1.5-2.5 meters. Skies were variable, but mostly clear.

Operations commenced at 0600 with a VMP deployment. An early morning MASCOT drop took place. Unsettled seas resulted in cancellation of small boat launch. MASCOT,Acoustic Buoy, RADCAM, and Hyperpro were deployed in the morning, a CTD (500 m, with bottles) at noon, and VMP, MASCOT + acoustic resonator, RADCAM and KLUDGE throughout the afternoon. VMP, KLUDGE, BIOLUME and MASCOT ops are planned for this evening and a 0100 CTD cast (w/o bottles) is planned. The wave LIDAR operated throughout.

All science equipment performed nominally, and all planned operations were successfully completed with the exception of those requiring small boat operations. The plan for tomorrow is attached.

Fantastic Labor Day BBQ on the sunny 02 deck - Filet mignon, grilled salmon, ranch beans, salad and corn on the cob! Much appreciated by all, and many thanks to the cook and stewards.

NOTE: Science meeting at 18:45.

Regards,

Marlon

Scientific Operations for 8 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/08/09

FLIP Status Report 070909

Note: Compiled at 0900 local time
Position: 17-32.141' N 157-40.776' W
Skies: Clear
Winds: 15 knots from 90 degrees
Seas: 3-4 ft from E direction
Swells: 4-6 ft from NNE direction
Pressure: 29.94 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 126 nm on an average course of 256 degrees
since going vertical. Currently we are on a straight westerl course. All
instruments for the starboard and port booms and the crowe's nest have now
been deployed and are running continuously. Ken Voss' group will be
deploying their DPOL and sky camera from the face boom and Marlon Lewis'
RADCAM from the starboard boom, Nick Hall Patch will be deploying his
acoustic resonator from the starboard side, and Ken Melville's group will
be deploying their underwater imaging/turbulence package from the port
boom. General operations meeting took place at 1700 last night and science
video meeting with the KM at 1900. Plans for deinstall will occur later
this week when we confirm tugboat details.

Nick

--------------------------------------------------------------------------------------09/07/09

6 September, 2009. KM drifting alongside FLIP, position at 20:00(ZT) was 17 deg 31.1'N 157 deg 29.7'W. Winds were ~ 15-20 knots throughout the day, with seas/swells ~ 1.5-2.0 meters. Skies were variable, but mostly clear.

Operations commenced at 0600 with a VMP deployment. An early morning MASCOT drop took place. The small boat was launched to deploy the AUV and then took Oliver and Ian for surface microlayer sampling. MASCOT, Acoustic Buoy, RADCAM, and Hyperpro were deployed in the morning, a CTD (500m, with bottles) at noon, and VMP, MASCOT and RADCAM throughout the afternoon. The acoustic resonator was successfully deployed as well with the new parts. The AUV was recovered in another small boat deployment. A 0100 CTD cast (w/o bottles) is planned. The wave LIDAR operated throughout.

All science equipment performed nominally, and all planned operations were successfully completed. The plan for tomorrow is attached.

NOTE: Science meeting at 18:45 after Labor Day bb

Regards,

Marlon

Scientific Operations for 7 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/07/09

FLIP Status Report 060909

Note: Compiled at 0800 local time
Position: 17-31.315' N 157-23.718' W
Skies: Clear
Winds: 15 knots from 80 degrees
Seas: 2-4 ft from E direction
Swells: 4-5 ft from NNe direction
Pressure: 29.90 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phones not operational
Comments: FLIP has now drifted 111 nm on an average course of 255 degrees since going vertical. We are turning in a more westerly direction from out original WSW course. Small boat transfer for plastic garbage and cable transfer to KM was successful yesterday morning. All instrumentation on the starboard boom has been deployed and is running continuously. An albedo sensor will be deployed for Marlon Lewis' group today on the port boom.s Otherwise all instrumentation has been installed on this boom as well and is running continuously. Yesterday Ken Melville's group deployed two additional cameras on the crowe's nest. Today Eric Terrill's WAMOS will be deployed from the upper deck. With respect to subsurface instruments, Ken Voss' group again deployed their sky camera and DPOL during the daylight hours from the face boom. Nick Hall Patch deployed his acoustic resonator float off the starboard side through the night. Ken Melville's group deployed their underwater imaging/turbulence package off the port side. Today Dariusz Stramski's group is scheduled to use the port side boom. Otherwise the schedule remains the same as yesterday for subsurface deployments.

Nick

--------------------------------------------------------------------------------------09/06/09

5 September, 2009. KM drifting alongside FLIP, position at 20:00(ZT) was 17 deg 31.4'N 157 deg 15.2'W. Winds were ~ 15-20 knots throughout the day, with seas/swells ~ 1.5-2.0 meters. Skies were variable, but mostly clear.

Operations commenced at 0600 with a VMP deployment. The small boat was deployed, and picked up parts from FLIP. It then deployed the AUV and took Oliver and Ian for surface microlayer sampling. MASCOT, Acoustic Buoy, RADCAM and HyperPro, were deployed in the morning, a deep CTD (2000 m, with bottles) at noon, and VMP and TSRB throughout the afternoon. The AUV was successfully recovered in another small boat deployment. Evening sampling includes VMP, Biolume and MASCOT; a 0100 CTD cast (w/o bottles) is planned as well. The wave LIDAR required a reset, but appeared to have operated all day.

All science equipment performed nominally. The acoustic resonator was repaired with the part from FLIP. All planned operations were successfully completed. The plan for tomorrow is attached.

NOTE: Science meeting at 18:45 and 19:00 videocon with FLIP crew.

Regards,

Marlon

Scientific Operations for 6 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/06/09

FLIP Status Report 050909

Note: Compiled at 0800 local time
Position: 17-32.814' N 157-05.875' W
Skies: Partially cloudy
Winds: 17 knots from 80 degrees
Seas: 3-5 ft from E direction
Swells: 4-6 ft from NNE direction
Pressure: 29.87 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phone not operational
Comments: FLIP has now drifted 94 nm on a course of 253 degrees since going vertical. Yesterday Ken Melville's group deployed a small camera and LED display to the port boom. All instrumentation for the starboard and port booms has now been deployed and will be running continuously. Melville's group will add two more cameras to the crowe's nest today. Other than that, all continuously measuring instrumentation has been installed. Yesterday Ken Voss' group deployed their sky camera and DPOL from the face boom, Nick Hall Patch deployed his acoustic resonator float off the starboard side, and Dariusz Stramski's group deployed their porcupine instrument from the port boom. Today the same subsurface instrument schedule as above will follow with the addition of an underwater turbulence/imaging package from the starboard boom by Ken Melville's group during the evening hours.

Nick

--------------------------------------------------------------------------------------09/05/09

4 September, 2009. KM drifting alongside FLIP, position at 20:00(ZT) was 17 deg 334.2'N 156 deg 56.5'W. Winds were ~ 20-25 knots throughout the day, with seas/swells ~ 2-2.5 meters. Skies were variable, but mostly clear.

Operations commenced at 0600 with a VMP deployment. High winds and seas again resulted in cancellation of small boat operations. MASCOT, RADCAM and HyperPro,were deployed in the morning, the CTD (500> m, with bottles)at noon, and MASCOT, Acoustic Buoy, VMP, RADCAM and KLUDGE throughout the afternoon. A 0100 CTD cast (w/o bottles) is planned. The wave LIDAR appeared to have operated all day. Full video communication was established with FLIP this evening thanks to Russel and Helen.

All science equipment performed nominally. The acoustic resonator requires a part currently on FLIP. All planned operations with the exception of those requiring the small boat were successfully completed.

The plan for tomorrow is attached.

NOTE: Science meeting at a time TBD with FLIP crew.

Regards,

Marlon

Scientific Operations for 5 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/05/09

3 September, 2009. KM drifting alongside FLIP, position at 21:30(ZT) was 17 deg 37.3'N 156 deg 41.2'W. Winds were ~ 20-25 knots throughout the day, with seas/swells ~ 2.5-3 meters. Skies were variable.

Operations commenced at 0600 with a VMP deployment. High winds and seas again resulted in cancellation of small boat operations. MASCOT, RADCAM and HyperPro,were deployed in the morning, the CTD (500> m, with bottles)at noon, and MASCOT, Acoustic Buoy, VMP, RADCAM and KLUDGE throughout the afternoon. VMP, Biolume, MASCOT and KLUDGE are planned for this evening and a 0100 CTD cast (w/o bottles) is planned. The wave LIDAR appeared to have operated all day.

All science equipment performed nominally. The acoustic resonator requires a part currently on FLIP. All planned operations with the exception of those requiring the small boat were successfully completed.

The plan for tomorrow is attached.

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 4 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/04/09

FLIP Status Report 040909

Note: Compiled at 0800 local time
Position: 17-36.593' N 156-49.274' W
Skies: Clear
Winds: 23 knots from 80 degrees
Seas: 4-6 ft from E direction
Swells: 5-7 ft from NNe direction
Pressure: 29.89 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phone not operational
Comments: Tom Golfinos is now able to access his emails. FLIP has now drifted 79 nm on a course of 250 degrees since going vertical. Yesterday Chris Zappa added his polarimeter system to the starboard boom. All instruments for this boom have now been deployed and are running continuously.
Ken Melville's group deployed their stereo IR system on the port boom. The LED display and a small camera will be deployed today and will be the final instruments deployed for the port boom. Ken Voss' group again deployed their sky camera and DPOL during the daylight hours from the face boom.
Nick Hall Patch deployed his acoustic resonator float off the starboard side. Dariusz Stramski's group deployed their optical porcupine instrument from the port side boom. Today the same subsurface instrument schedule as above will follow with the addition of an underwater turbulence/imaging system from the starboard boom from Ken Melville's group added during the evening hours.

Nick

--------------------------------------------------------------------------------------09/04/09

2 September, 2009. KM drifting alongside FLIP, position at 19:00(ZT) was 17 deg 39.6'N 156 deg 24.5'W. Winds were ~ 20-25 knots throughout the day, with seas/swells ~ 2.5-3 meters. Skies were variable.

Operations commenced at 0600 with a VMP deployment. High winds and seas resulted in cancellation of small boat operations. MASCOT, RADCAM and HyperPro,were deployed in the morning, the CTD (500> m, with bottles)at noon, and MASCOT, VMP and RADCAM throughout the afternoon. A 0100 CTD cast (w/o bottles) is planned. The wave LIDAR appeared to have operated all day.

All science equipment performed nominally. The acoustic resonator is being evaluated. All planned operations with the exception of small boat operations were successfully completed.

The plan for tomorrow is attached.

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 3 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/03/09

FLIP Status Report 030909

Note: Compiled at 0930 local time
Position: 17-39.113' N 156-33.613' W
Skies: Clear
Winds: 22 knots from 85 degrees
Seas: 4-6 ft from E direction
Swells: 6-8 ft from NNE direction
Pressure: 29.89 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phone not operational
Comments: Currently we have drifted 64 nm with a course of 250 degrees since going vertical. Installation of some scientific equipment will continue today along with vigorous sampling. Yesterday Chris Zappa added his imaging system on the starboard boom. He will add his polarimeter imaging system today, ending installations for the starboard boom. Ken Melville's group added both their stereo visible and stereo IR systems to the port side boom. Installation of the LED display for their underwater imaging system and crowe's nest cameras is scheduled for today, ending installations for the port side boom. Dariusz Stramski's group deployed their porcupine instrument from the port boom for the majority of the day yesterday. Ken Voss' group deployed their sky camera and DPOL again from the face boom. These underwater deployments are scheduled to continue today.

Nick

--------------------------------------------------------------------------------------09/03/09

1 September, 2009. KM drifting alongside FLIP, position at 19:00(ZT) was 17 deg 43.3'N 156 deg 10.83'W. Winds were ~ 18-20 knots throughout the day, with seas/swells ~ 1.5-2 meters. It was cloudy most of the day.

Operations commenced at 0600 with a VMP deployment. The small boat was launched, and successfully delivered Russel to the FLIP, and picked up overflow gash. The boat returned to KM and picked up Ian, Oliver and the AUV, which was successfully launched. MASCOT and HyperPro,were deployed in the morning, the CTD (500> m, with bottles)at noon, and MASCOT, VMP and RADCAM throughout the afternoon. The AUV was recovered at 1500, and collected good data. VMP, Biolume and LISST deployments are scheduled for this evening and a 0100 CTD cast (w/o bottles) is planned for 0100. The wave LIDAR operated all day

All science equipment performed nominally, and all planned operations were successfully completed

The plan for tomorrow is attached.

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 2 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/02/09

FLIP Status Report 020909

Note: Compiled at 1000 local time
Position: 17-41.493' N 156-19.245' W
Skies: Cloudy
Winds: 22 knots from 80 degrees
Seas: 4-6 ft from E direction
Swells: 7-9 ft from N direction
Pressure: 29.87 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 85 gallons
Communications: Cell phone not operational
Comments: Yesterday all the winches and air tuggers were setup on the lower deck for science usage. Science party continues deployment of instrumentation today. Yesterday Chris Zappa deployed his eddy flux system and Riegl altimeters, Russel Morison deployed his scanning LIDAR system, and Johannes Gemmrich deployed his visible cameras all along the starboard boom. Ken Melville's group deployed their Cadillac system, thermistor chain, and scanning LIDAR along the port side boom. All these instruments will now run continuously throughout the experiment. Ken Voss' group deployed both their sky camera and DPOL for the majority of the day off the face boom. These cameras will again be deployed this afternoon along with the Porcupine instrument from Dariusz Stramski's group. Science party meeting again scheduled for 1900 this evening. Note: we have now drifted 50 nm at a course of 250 degrees since going vertical.

Nick

--------------------------------------------------------------------------------------09/02/09

UCSB R/P Flip Radyo Sensors

--------------------------------------------------------------------------------------09/01/09

FLIP Status Report 010909

Note: Compiled at 0900 local time
Position: 17-46.257' N 156-04.490' W
Skies: Cloudy
Winds: 18 knots from 70 degrees
Seas: 3-4 ft from E direction
Swells: 5-6 ft from NNE direction
Pressure: 29.88 in and rising
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phone not operational
Comments: WAMOS mast from Eric Terrill's group deployed yesterday by FLIP
crew. Successful boat transfer of Russel Morison this morning at 0800 from
Kilo Moana. Ken Voss' group deployed their DPOL and sky camera in the
afternoon hours yesterday. Ken Melville's group installed their eddy flux
system and altimeter for sampling yesterday as well. Installation and
deployment of science equipment continues today. Science meeting at 1900
this evening to discuss sampling strategy and boom usage.

Nick

--------------------------------------------------------------------------------------09/01/09

31 August, 2009. KM drifting alongside FLIP, position at 19:30(ZT) was 17 deg 48.3'N 155 deg 59.3'W. Winds were ~ 20 knots most of the day, with seas/swells ~ 2+ meters. Overcast conditions and squalls were prevalent in the morning, with clear skies in the afternoon. FLIP is setting more towards the south now (~240 T), and averaging 0.6 knots of drift.

Operations commenced at 0600 with a VMP launch. Small boat operations were cancelled due to sea conditions and high winds. MASCOT, RADCAM, HyperPro,were all deployed in the morning, the CTD (500> m, with bottles) at noon, and Secchi disk (26 meters) MASCOT, VMP and acoustic buoy throughout the afternoon. The AUV launch was cancelled, again due to high seas. A 0100 CTD cast (w/o bottles) is planned for this evening. The wind LIDAR operated all day; it was removed for cleaning in the late afternoon and reinstalled.

All science equipment performed nominally, and all planned operations were successfully completed with the exception of those using the small boat. Repositioning of the ship following KM operations to lie alongside FLIP occupied time throughout the day.

The plan for tomorrow is attached. We will attempt small boat operations, weather permitting (including personnel transfer, microlayer sampling, and AUV deployment/recovery).

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 1 Sept., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------09/01/09

Marlon

Scientific Operations for 31 August., R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------08/31/09

FLIP Status Report 310809

Note: Compiled at 0800 local time
Position: 17-50.488' N 155-52.953' W
Skies: Cloudy with rain
Winds: 25 knots from 110 degrees
Seas: 4-5 ft from E direction
Swells: 5-7 ft from NE direction
Pressure: 29.87 in and rising
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 87 gallons
Communications: Cell phones not operational
Comments: All booms were deployed successfully yesterday under clear and windy conditions. Scientific equipment installation has begun and will continue today. Due to weather conditions the scheduled boat transfer for personnel exchange has been canceled today. Currently FLIP is drifting on a course of 260 degrees at 0.6 knots. Since disconnecting from the Navajo we have drifted a total of 24 nm.

Nick

--------------------------------------------------------------------------------------08/31/09

FLIP Status Report 300809

Note: Compiled at 0730 local time
Position: 17-56.034' N 155-40.931' W
Skies: Cloudy
Winds: 15 knots from 60 degrees
Seas: 3-5 ft from E direction
Swells: 4-5 ft from SE direction
Pressure: 29.89 in and rising
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 80 gallons
Communications: Cell phone not operational
Comments: FLIP reached initial starting position via the USNS Navajo at at local time of 1400 on 290809 where it disconnected its tow line. At 1700 hours FLIP went vertical and started drifting. The position at that hour was 17-58.260' N 155-34.249' W. Currently we are drifting on a course of 250 degrees at a speed of 0.5 knots. Today the crew is preparing to deploy the booms. The space in the science labs has been allotted and the organization of science party equipment has begun. Once the booms are deployed, instrument installation will begin. Science party meeting scheduled for 1900.

Nick

--------------------------------------------------------------------------------------08/30/09

29 August, 2009. KM on station, 18o 0.0' N, 155o 30W (most of day). Winds were ~ 20 knots all day, with seas/swells ~1.5 meters. It was primarily cloudy with rain squalls, interspersed with clear periods.

We arrived back at the high seas insertion point this morning at 0600 in time for a VMP launch, with the new (and working) termination. MASCOT, RADCAM, HyperPro, and TSRB were all deployed in the morning, the CTD (500 m, with bottles) at noon, and MASCOT, VMP, KLUDGE and acoustic buoy throughout the afternoon. The VMP and BIOLUME will deploy this evening prior to an 0100 CTD cast (w/o bottles). The wind LIDAR operated all day, and the ship-to-ship comms link was successfully tested internally.

All science equipment performed nominally, and all planned operations were successfully completed.

FLIP was towed alongside at ~1400, and after release from the tow vessel, drifted downwind for 5 miles and then underwent successful rotation. KM launched a small boat to FLIP and picked up spares for the AUV. FLIP continued to drift at ~1 knot after it was vertical, which is 2-3 times what I believe we had estimated for a drift rate. We are currently (20:30 local) 2 M south, and 6 M west of the insertion point. We will be challenged to remain close to FLIP tomorrow so plan several repositioning periods (see attached).

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 28 August, R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------08/30/09

Hello all,

FLIP met the Kilo Moana at the previously defined initial posit of 18-0.0' N 155-30' W at 1400 today via USNS Navajo. FLIP went into vertical position at 1700 under 20 knot winds and 3-5 ft seas. Currently FLIP is drifing with a course of 260 degrees at a speed of 0.7 knots. As of 2200 local time our location is 17-57.879' N 155-37.604' W. Full vitals reports will continue tomorrow morning.

Regards,
Nick

--------------------------------------------------------------------------------------08/30/09

28 August, 2009. KM on station, 19o 15.0N, 155o 58W. We spent today within sight of Moana Loa, 4 M offshore in 2000 m of water. Weather was calm, with glassy surface throughout most of the day, albeit with light swell. It was very warm, with hazy sun (much VOG).

Operations commenced at 0745 with the launch of the small boat and like the day before, the acoustic buoy and MASCOT were deployed during this sampling, and recovered on return. Trichodesmium remained, although not as dense as the previous station. RADCAM was deployed, followed by a (500 m) CTD at noon with water samples. The small boat redeployed at 1300. The KLUDGE was launched and recovered, followed by MASCOT with acoustic resonator. After recovery of the small boat, Hyperpro and RADCAM were successfully launched, followed by the vertical microstructure profiler (VMP).

All planned deployments were therefore successfully completed, with few technical difficulties. We are working a ship-to-ship comms link, and resolved minor issues with the CTD sheave and VMP termination.

The plan for tomorrow is attached. Note that we are currently enroute back to the High Seas site with an ETA of 0600. We understand that FLIP is enroute as well, and will join us mid-day.

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 28 August, R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------08/29/09

FLIP Status Report 290809

Note: Compiled at 0900 local time
Position: 18-26.486' N 155-50.786' W
Skies: Cloudy
Winds: 25 knots from 160 degrees
Seas: 5-6 ft from SE direction
Swells: 6-7 ft from SE direction
Pressure: 29.89 in and steady
Equipment: All FLIP equipment operational
Fuel consumed last 24 hours: 50 gallons
Communications: Cell phone not operational
Comments: Currently in transit to location via USNS Navajo. Traveling at
8.8 knots with a heading of 137 degrees. Plan to be on site at 1300.
Depending on the weather conditions that we find on location, we'll make
the decision whether to flip vertical. Science reports will commence once
in the vertical position.

Nick

--------------------------------------------------------------------------------------08/29/09

http://www.noisemakers.org.uk/modules/blogs/blog.cfm?id=5

--------------------------------------------------------------------------------------08/28/09

27 August, 2009. KM on station, 19o 15.0N, 156o 10W. We arrived at this station at 07:30 after steaming through somewhat rough seas. In the lee however, winds were <10 knots, and seas were small with 3' swell. Skies ranged from partly cloudy to clear throughout most of the day, although very hazy.

Operations commenced at 0745 with the launch of the small boat (zodiac) with Oliver and Scott P. onboard for microlayer sampling. While they were out, the acoustic buoy and MASCOT were deployed, and recovered on return of the small boat at 10:00. The surface waters were covered with Trichodesmium which created multiple surface slicks in all directions (sampled by Oliver, Dariusz and Pierre). RADCAM and HyperPro were then deployed, followed by a deep (1000 m) CTD at noon with water samples Vertical structure was consistent with previous and Secchi depth was 27 meters. The small boat redeployed at 1300 with Ian and Oliver, and again, MASCOT was redeployed with acoustic resonator. After recovery of the small boat, KLUDGE2 was successfully launched. Tonight, microstructure and bioluminesence will be measured by Scott P and Hemantha.

All planned deployments were therefore successfully completed, with few technical difficulties.

The plan for tomorrow is attached. Note that the ship has now repositioned this evening to a site 4 M offshore (19 15'N 155 58'W) for work tomorrow. We plan to return to the High Seas site departing early tomorrow evening.

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 28 August, R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------08/28/09

Satellite images

hi_day257_264

hi_night257_264

hi_chlo_modis257_264

hi_day249_256

hi_night249_256

hi_chlo_modis249_256

hi_day241_248

hi_night241_248

hi_chlo_modis241_248

hi_day233_240

hi_night233_240

hi_chlo_modis233_240

hi_day225_232

hi_night225_232

hi_chlo_modis225_232

hi_chlo_seawifs225_232

hi_day217_224

hi_night217_224

hi_chlo_modis217_224

hi_chlo_seawifs217_224

hi_day209_216

hi_night209_216

hi_chlo_modis209_216

hi_chlo_seawifs209_216

--------------------------------------------------------------------------------------08/28/09

26 August, 2009. KM on station, 18o 0.0?N, 155o 30?W. Winds increased to 20-25, gusting to 30 knots during squalls, and with sea/swell 2.5-3 meters. Skies ranged from partly cloudy to overcast throughout most of the day, with rain occasionally in squalls.

Operations commenced 0100 with a CTD cast to 300 meters (no bottles) At 0530 the tethered free-fall microstructure profiler was deployed followed sequentially by multiple MASCOT, RADCAM and HyperPro deployments. The CTD was deployed to 500 meters at noon and water samples were taken from multiple depths. Vertical structure was consistent with previous ? mixed layers 45-50 meters, chlorophyll maximum at 120 meters, 1% light also at ~120 meters. Secchi depth was 25 meters. Microstructure was evaluated again and MASCOT redeployed with acoustic resonator, followed by KLUDGE2. An Acoustic buoy was put together and deployed successfully (Helen, Oliver). The LIDAR system ran successfully throughout the day. A last series of microstructure profiles will take place shortly.

All planned deployments were therefore successfully completed, with few technical difficulties.

The plan for tomorrow is attached. Note that the ship will be repositioned this evening into the lee of the island of Hawaii and Thursday will be spent at a site 15 M off the coast, and Friday at a site 4 M offshore. We plan to return to the High Seas site for Saturday morning.

NOTE: Science meeting at 18:45 tomorrow following celestial event.

Regards,

Marlon

Scientific Operations for 27 August, R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------08/27/09

Scientific Operations, R/V Kilo Moana RaDyO

--------------------------------------------------------------------------------------08/25/09

RaDyO Hawaii Forecasts

The forecasts are available at https://cavu.nrlmry.navy.mil/COAMPSOS/

click on the "Hawaii" AREA link. The project will be updated twice a day.

A map of the forecast area is attached; COAMPS is running for all three nests and NCOM is running for the inner two nests; SWAN is running for the outermost nest and will (hopefully) run for the innermost later today.

The default page is the COAMPS IPVS graphics, one matrix of products by forecast time per nest. Click "GMT Products" tab near the top of the page to see a similar display that includes additional COAMPS products and all the SWAN and NCOM products. Look out for Hilda. The SWAN wave forecast is impressive.

See also:

http://www.nrlmry.navy.mil/tc-bin/tc_home2.cgi?YEAR=2009&MO=08&BASIN=EP AC&STORM_NAME=11E.HILDA&PROD=track_vis&PHOT=yes&ARCHIVE=active&NAV=tc&AG E=Latest&SIZE=Thumb&STYLE=tables&AID_DIR=/SATPRODUCTS/kauai_data/www/EPA C/tropics/track_vis/dmsp&TYPE=ssmi/

John

--------------------------------------------------------------------------------------08/25/09

All,

attached is a map showing the location of the pier where FLIP will berthed in HI (V6 Pier). Also attached are directions from the Hilton Hawaiian Village Hotel.

Direction to Flip
Luc

--------------------------------------------------------------------------------------08/11/09

hi Tommy,

I ordered some Envisat wide swath data for experiment. I am not used to lower latitude coverage, so I was surprised there were so few options really, just not much coverage. The dates are Sept 1, 3,4, 6, 7, 9, 10. They should cover Flip and big Hawaii or other islands abit. Now on to images from Sept.

cheers, Ben

--------------------------------------------------------------------------------------08/04/09

Tommy et al.,

A few years ago when I submitted a NSF proposal to study the Hawaii eddies, I have studied the surface wind from QuikSCAT, the blended wind as mentioned, and a few other alternative sources.

The best wind around the Hawaii Islands is from a mesoscale atmospheric model. There is a faculty in the Department of Meteorology, University of Hawaii at Manoa, who has done extensive research in the atmospheric and wind for this region. His contact information is below, and a web link to his real-time WRF forecast at 6-km resolution. He has a number of WRF domains with higher resolutions, but over a smaller area.

Prof. Yi-Leng Chen
yileng@hawaii.edu

http://www.soest.hawaii.edu/MET/Faculty/wrf/arw/arw_state_6km.html

Let me know if you have any questions.
Regards. -Yi

--------------------------------------------------------------------------------------07/30/09

Oliver and Svein have a paper in press in Marine Chemistry using their RaDyO data sets. The paper is available in the Publications section of this website.

Also, their work is described in the following NY Times article which shows a picture of KM and Lil KM:

http://www.nytimes.com/2009/07/28/science/28ocea.html?_r=1&hpw

--------------------------------------------------------------------------------------07/30/09

SST website from Yi Chao:

http://sst.jpl.nasa.gov/SST

--------------------------------------------------------------------------------------07/30/09

Leila Carvalho has found a really nice set of weather maps relevant for the SBC experiment. The site is:

http://squall.sfsu.edu/crws/archive/wcsathts_arch.html

The modeling work that she is doing with collaborators will give us smaller scale view of the weather over the SBC during our experiment.

--------------------------------------------------------------------------------------07/29/09

Lil KM makes the NY Times!!! Go to the following website to see Lil KM picture and story:

Science Link

--------------------------------------------------------------------------------------07/28/09

All,
fyi, since August 21st is a Hawaiian state holiday, crane services will only be available on August 20th to load science equipment on FLIP. (only hand carried items will be loaded on the 21st). Does anyone expect extensive use of the crane for additional operations (e.g. mounting heavy equipment on FLIP's hull etc...)?
Please let me know asap.
Luc

--------------------------------------------------------------------------------------07/23/09

RaDyO'ers,
the new VSAT internet telecommunication system was successfully tested during the HiRes FLIP Experiment a couple of weeks ago. As of now, the bandwidth for the RaDyO HI cruise will be 64kbps/64kbps (ship to shore/shore to ship) For an additional $1.5k/mo we would be able to get 256kbps/256kbps, for $2.5/mo we would get a bandwidth up to 512kbps/1Mbps. Anyone interested in increasing the bandwidth? We could spread the cost among FLIP PIs if that's the case.
Let me know.
Luc

--------------------------------------------------------------------------------------07/23/09

Steve,
The new sewage tanks and port boom were tested successfully last week during the HiRes FLIP pilot. We also recorded very promising results from the Stereo IR system that we will deploy in HI, we will present preliminary results at the Symposium.
Luc

--------------------------------------------------------------------------------------07/20/09

All,
below is an updated and final schedule for FLIP operations in Hawaii.
*We will NOT depart on August 23rd, but on the 24th instead due to
limited harbor tug availability on Sunday.
*

FLIP 2009 Hawaii Schedule
03 August, 0800, Sioux moor north side Nimitz pier and load FLIP mooring
for transport to PH.
03 August, about 1300, Sioux depart for PH with FLIP in tow.
07 August, Gaines travel to Hawaii
14 August, about 1400 Sioux/FLIP arrive PH. Submit stores order to FISC.
17 August, AM offload mooring gear from Sioux at the FISC pier. The
mooring gear will be stored on the pier and nearby. FLIP crew assist
with the offload. RaDyo group arrive Hawaii.
20/21 August, load FLIP, stores delivery.
*24 August, 0800, Navajo U/W with FLIP and transit to an area south of
the Island of Hawaii, vicinity18N, 155-30W.
2*5 August, FLIP commence drift cruise, Navajo return to PH 26 August
15 Sept Navajo, recover FLIP and tow to PH, expect PM 16 or AM 17 Sept
RTP

Unloading is schedule for September 17-18 (depends on arrival date)


Best,
Luc
--


Luc Lenain wrote:
All,
few updates for the upcoming FLIP experiment.
Bill Gaines was in Hawaii two weeks ago to prepare FLIP's arrival in
Pearl City in August.
*
*Final FLIP schedule is as follow:*

03-14 Aug 09, tow SDGO to PH, load mooring for transport only and U/W 03
Aug, arrive PH and offload 14 Aug.

23-24 Aug 09, ONR RaDyo Salvor tow FLIP to drift location, detach FLIP,
Salvor RTP.

14-16 September 09, Salvor recover FLIP and RTP PH.

28-30 September 09, ONR/N45 mammal cruise, Salvor load mooring AM 28
Sept U/W tow FLIP to mooring area, 29 Sept moor FLIP, Salvor RTP.*

* (Salvor is the tug, PH:Pearl Harbor)*

*Shipping:* We are still finalizing the shipping arrangements to Pearl
Harbor - I will know more in a couple of days (shipping address etc...),
and will pass the information right away. Please let me know if you are
planning on loading equipment on FLIP in San Diego. This would require
coordination with the Shop/FLIP's crew ahead of time, in particular if
the items are heavy and require a crane.

*Loading:* Loading dates in Pearl City are set for August 20&21. A crane
will be available on both days.

*Access to the Base:* Bill is still finalizing arrangements with SPAWAR
- I need a list of names of all personnel who will have to get on the
base/FLIP. This list has to be finalized by the end of the month. Please
let me know who will need to access the base for each of your respective
groups. At this point they are not requiring TWIC cards. However it
would be best for all US personnel to apply for it anyway.

*Internet:* Through the HiRes DRI program, we were able to purchase a
VSAT communication package for FLIP (KVH V7) - The system will be tested
for the first time during the HiRes FLIP pilot experiment in July 09.
Assuming the initial deployment is successful, we will be able to use
this system in Hawaii, with slowest bandwidth configuration (64kbps
monthly plan - unlimited) - at _no costs_ for the PIs. If anyone is
interested, there is the possibility of increasing the bandwidth, e.g.
increasing it to 512kbps ship to shore / 1/024kbps shore to ship would
cost $3.5k for one month. I can provide more details if needed.

*Port boom:* the new port boom was successfully tested in early May 2009.

*Starboard boom:* We are getting steps built to ease the access to
starboard boom.

*Sewage Tanks:* Two 500 gallons sewage tanks are being installed on FLIP
and will be tested in July during the HiRes Pilot. We will still need to
discharge the sewage on a daily basis, but this will provide us with
control, and set the time of discharge during a low sampling period.

*Air Conditioning:* FLIP crew has been working on a complete new set of
A/C units - Again this will be tested during the HiRes pilot in July 09.

*Lab space:* Following discussions at the last meeting, I am not
planning on taking the blue rack that Mirek and Russell used. Please let
me know if you do want to use it.

Please let me know if you have additional questions/concerns.
Best,
Luc


Luc Lenain wrote:

Hi All,
it was nice to see you guys.
FLIP will be moored in Pearl Harbor at V5 or V6 (see attachment (sorry
for the quality) and
http://maps.google.com/maps?hl=en&client=firefox-a&q=hawaii+pearl+city+
map&ie=UTF8&split=0&ll=21.3757,-157.969408&spn=0.017664,0.033045&t=h&z=
16 )
We will have covered space secured for storing shipment at the nearby
SPAWAR building.
Please let me know if you have specific needs (crane, staging space
....) while at dock.
Derek, we will have access to a crane to install your optical packages
on FLIP.
Luc
--

--------------------------------------------------------------------------------------06/22/09

All,

The address for the shipment to Hawaii is:

Attention to:

Bernie Donathan
MSCO Pearl Harbor
(RaDyO R/P FLIP Experiment)
2293 Victor Wharf Access Road
Pearl City, HI. 96782
808-471-4564

Please consolidate all gear in one shipment per group, they have very limited staff on location.

Access to the base: I just received their final requirements. No need for TWIC card, but I need names, nationality and organization affiliation to gain access to the pier and FLIP asap.

Best,

Luc

--

--------------------------------------------------------------------------------------06/15/09

All,

few updates for the upcoming FLIP experiment.

Bill Gaines was in Hawaii two weeks ago to prepare FLIP's arrival in Pearl City in August.

*

*Final FLIP schedule is as follow:*

03-14 Aug 09, tow SDGO to PH, load mooring for transport only and U/W 03

Aug, arrive PH and offload 14 Aug.

23-24 Aug 09, ONR RaDyo Salvor tow FLIP to drift location, detach FLIP, Salvor RTP.

14-16 September 09, Salvor recover FLIP and RTP PH.

28-30 September 09, ONR/N45 mammal cruise, Salvor load mooring AM 28 Sept U/W tow FLIP to mooring area, 29 Sept moor FLIP, Salvor RTP.*

* (Salvor is the tug, PH:Pearl Harbor)*

*Shipping:* We are still finalizing the shipping arrangements to Pearl Harbor - I will know more in a couple of days (shipping address etc...), and will pass the information right away. Please let me know if you are planning on loading equipment on FLIP in San Diego. This would require coordination with the Shop/FLIP's crew ahead of time, in particular if the items are heavy and require a crane.

*Loading:* Loading dates in Pearl City are set for August 20&21. A crane will be available on both days.

*Access to the Base:* Bill is still finalizing arrangements with SPAWAR - I need a list of names of all personnel who will have to get on the base/FLIP. This list has to be finalized by the end of the month. Please let me know who will need to access the base for each of your respective groups. At this point they are not requiring TWIC cards. However it would be best for all US personnel to apply for it anyway.

*Internet:* Through the HiRes DRI program, we were able to purchase a VSAT communication package for FLIP (KVH V7) - The system will be tested for the first time during the HiRes FLIP pilot experiment in July 09. Assuming the initial deployment is successful, we will be able to use this system in Hawaii, with slowest bandwidth configuration (64kbps monthly plan - unlimited) - at _no costs_ for the PIs. If anyone is interested, there is the possibility of increasing the bandwidth, e.g. increasing it to 512kbps ship to shore / 1/024kbps shore to ship would cost $3.5k for one month. I can provide more details if needed.

*Port boom:* the new port boom was successfully tested in early May 2009.

*Starboard boom:* We are getting steps built to ease the access to starboard boom.

*Sewage Tanks:* Two 500 gallons sewage tanks are being installed on FLIP and will be tested in July during the HiRes Pilot. We will still need to discharge the sewage on a daily basis, but this will provide us with control, and set the time of discharge during a low sampling period.

*Air Conditioning:* FLIP crew has been working on a complete new set of A/C units - Again this will be tested during the HiRes pilot in July 09.

*Lab space:* Following discussions at the last meeting, I am not planning on taking the blue rack that Mirek and Russell used. Please let me know if you do want to use it.

Please let me know if you have additional questions/concerns.

Best,

Luc

--------------------------------------------------------------------------------------06/15/09

Hi All,

it was nice to see you guys.

FLIP will be moored in Pearl Harbor at V5 or V6 (see attachment (sorry for the quality) and http://maps.google.com/maps?hl=en&client=firefox-a&q=hawaii+pearl+city+m ap&ie=UTF8&split=0&ll=21.3757,-157.969408&spn=0.017664,0.033045&t=h&z=16 ) We will have covered space secured for storing shipment at the nearby SPAWAR building.

Please let me know if you have specific needs (crane, staging space ....) while at dock.

Derek, we will have access to a crane to install your optical packages on FLIP.

Luc

--

--------------------------------------------------------------------------------------06/15/09

Hi Tommy,

based on the existing current and wind climatology, we picked 18:00 N, 155.5 W, as a start posit for FLIP's drift. (about 28hrs steam from PH for the tug). FLIP drift with the current, and orient itself in the wind, i.e. FLIP's hull facing the upcoming wind.
Luc --

--------------------------------------------------------------------------------------05/26/09

Below are two informational pieces from K/M Marine Tech Elly Speicher:

1) Deck Layouts (PDF file)

2) UH Network Access Control Policy (PDF file)

--------------------------------------------------------------------------------------05/12/09

April 9, 2009 Email updates on FLIP dumping and nightime sampling from KM and FLIP

Tommy,

FLIP will be equipped with a 1,000 gallons holding tank (2x 500gal tanks). it will be purged everyday, tentatively in the middle of the night. (we will generate ~750gal sewage /day, including sea water used in the flushing process...)

Luc

--------------------------------------------------------------------------------------04/09/09

Email from Gray Drewry (Dec. 4, 2008) of University of Hawaii Ship operations concerning shipping instructions for Kilo Moana:

Please copy myself, Keith and SNUG on e-mails concerning shipping packages to our facility, we appreciate a heads up. What is most importanted is the correct labling of packages
All deliveries should be clearly marked as follows:

PI TOMMY DICKEY
R/V KILO MOANA RaDyO
C/O UH MARINE CENTER
#1 SAND ISLAND ACCESS ROAD
HONOLULU, HI 96819

If you have a container please go to the University of Hawaii Marine Center web page
http://www.soest.hawaii.edu/UMC/marinecenter.html
and down load a copy of the Container Authorization for Arrival Form to your computer, complete the form and attach it to an e-mail to myself (gray@soest.hawaii.edu), we will need this form prior nto any containers being allowed entry to our facility.
Contact Information:

University of Hawaii Marine Center
#1 Sand Island Access Road, Honolulu, HI 96819
Phone (808) 842-9813, Fax (808) 842-9833 Toll Free 1-888-800-0460

Stan Winslow, Marine Superintendent (808) 842-9814 email swinslow@soest.hawaii.edu
Gen Pickering, Administrative Assistant (808) 842-9810 email genp@soest.hawaii.edu
Gray Drewry , Port Operations Manager (808) 842-9815 gray@soest.hawaii.edu
John Nikola, Port Engineer (808) 842-9816 jnikola@soest.hawaii.edu
Keith Adams, Purchasing Agent (808) 842-9811 keitha@soest.hawaii.edu
UH Marine Center snug@soest.hawaii.edu

Phone numbers for Kilo Moana while in port (808) 842-9817 / 842-9834

At this time there are no security requirements other than your science team having proper government issued PHOTO ID, drivers Lic, Passport will suffice, also as your cruise gets closer please provide a list of participants

If you have any more questions please don't hesitate to ask
Regards
Gray

--------------------------------------------------------------------------------------02/20/09

Hi All,
It was nice to see you guys.
FLIP will be moored in Pearl Harbor at V5 or V6 (see attachment (sorry for the quality) and http://maps.google.com/maps?hl=en&client=firefox-a&q=hawaii+pearl+city+map&ie=UTF8&split=0&ll=21.3757,-157.969408&spn=0.017664,0.033045&t=h&z=16 )
We will have covered space secured for storing shipment at the nearby SPAWAR building.
Please let me know if you have specific needs (crane, staging space ....) while at dock.
Derek, we will have access to a crane to install your optical packages on FLIP.
Luc
--

--------------------------------------------------------------------------------------02/20/09

RaDyOer's,

Could you please let me know who from your respective groups will be participating in the Hawaii experiment FLIP component (name - email - phone #). I need to formalize the science party name list. Reminder -FLIP Schedule
03-14 Aug 09, Sioux tow SDGO to PH, load mooring for transport only and U/W

03 Aug, arrive PH and offload 14 Aug.
21-22 Aug 09 Loading
23-24 Aug 09, ONR RaDyo Salvor tow FLIP to drift location, detach FLIP, Salvor RTP.
14-16 September 09, Salvor recover FLIP and RTP PH. (Salvor will be operating in the Hawaii area for the duration of the experiment and will be available in the event of an emergency recovery)

Best,
Luc

--------------------------------------------------------------------------------------02/11/09